# GraphPilot: Grounded Scene Graph Conditioning for Language-Based Autonomous Driving

### v1

@@ 4.5 Training
When a scene graph is used during training (SG11 or SG10), it is integrated based on the selected abstraction level, serialization format, and injection template. These choices govern the form and placement of the graph in the prompt but do not alter the model architecture or optimization procedure. The total training time per run depends primarily on the abstraction level and serialization format, ranging from 7.5 hours (Actor-Only, Text) to 13 hours (Full, JSON). Among the formats, Text is the most efficient, followed by YAML, while JSON incurs the highest overhead due to its more verbose structure. Injection template has negligible effect on training time.
@@ 4.6 Results
Baselines. We first establish baselines by using the official implementations and pretrained checkpoints of LMDrive and BEVDriver, referred to as SG00-PT. These results serve as primary reference points to ensure fair, environment-matched comparisons. [^1] We report reproduced results rather than official numbers, as the authors of LMDrive acknowledge that performance varies significantly depending on GPU model as reported by other authors as well (see issues 37, 52, and 56 in LMDrive repository; links omitted for review according to author guidelines). Same behaviour was observed for BEVDriver. Additionally, to isolate the effect of scene graph supervision itself, we fine-tune both models on our dataset without scene graph input at training or test-time (SG00-FT). Table 1 shows that SG00 fine-tuning shifts performance in a model-specific way: LMDrive gains in safety (higher IS) but sees a drop in RC, while BEVDriver substantially improves RC and resulting in a higher DS. These outcomes provide a neutral reference for assessing the added value of scene graph conditioning.
@@
Specifically, we evaluate all combinations of abstraction level, serialization format, and injection template. Table 2 reports DS, RC, and IS for LMDrive averaged over the LangAuto Tiny, Short, and Long tracks. We give mean rows and columns to support comparisons across serialization formats and injection templates, respectively. Notably, multiple test-time scene graph configurations already surpass the reproduced baseline without scene graphs (Table 1), underscoring the potential of this approach even without additional training. Two consistent trends emerge: First, injection template V3 outperforms both V2 and V1, with V1 trailing significantly. Second, Actor-Only abstractions yield the best results across all formats, indicating that minimal yet structured relational information is most beneficial when applied to pretrained models.
These results highlight the strength of lean representations and well-designed prompts, even without additional training. Actor-Only graphs in particular strike a favorable balance between semantic relevance and token budget, introducing the fewest additional tokens across all formats, as shown in Table 3. Based on these insights, we select the Actor-Only abstraction and injection templates V2 and V3 for all subsequent fine-tuning experiments, focusing on configurations that maximize prompt efficiency and clarity.
@@
Configuration LMDrive BEVDriver
@@
Scene Graph-Conditioned Fine-Tuning. Table 4 reports results for SG11 and SG10 configurations, using the best-performing abstraction (Actor-Only) and injection templates (V2 and V3). For LMDrive, SG10 consistently outperforms SG11 in DS, and V3 surpasses V2, suggesting that exposure to structured context during training enhances internal planning behavior, even when that structure is absent at test-time. The best configuration (Text, Actor-Only, V3, SG10) achieves a DS of 51.8, marking a substantial improvement over the reproduced baseline (42.95) and SG00 fine-tuning (41.55). Moreover, SG10 and SG11 significantly outperform SG01 (44.7), demonstrating that scene graph-conditioned training is more effective than test-time injection only. BEVDriver shows complementary preferences: YAML slightly outperforms Text, and V2 is favored over V3. Nevertheless, SG10 again matches or exceeds SG11, with the top configuration (YAML, Actor-Only, V2, SG10) reaching a DS of 56.1, significantly outperforming both the reproduced baseline (44.70) and SG00 (48.94). We attribute differences in optimal serialization and templates in part to the underlying language backbone (LLaVA-v1.5 in LMDrive vs. LLaMA-7B in BEVDriver), but the broader pattern holds across both: SG10 provides consistent benefits without requiring graph input at test-time.
@@
Extended Infraction Analysis. To better understand how scene graph conditioning influences safety and behavior, we analyze detailed sub-metrics in Table 5. Rather than focusing on best-case results, we report averages over all V3-based LMDrive configurations (Text, JSON, YAML) to assess the generality of the observed effects.

### v2

@@ 4.5 Training
When a scene graph is used during training (SG11 or SG10), it is integrated based on the selected abstraction level, serialization format, and injection template. These choices govern the form and placement of the graph in the prompt but do not alter the model architecture or optimization procedure. The total training time (GPU hours) per run depends primarily on the abstraction level and serialization format, ranging from 61 hours (Actor-Only, Text) to 107 hours (Full, JSON) for LMDrive. Among the formats, Text is the most efficient, followed by YAML, while JSON incurs the highest overhead due to its more verbose structure. Injection template has negligible effect on training time.
@@ 4.6 Results
Baselines. We first establish baselines by using the official implementations and pretrained checkpoints of LMDrive and BEVDriver, referred to as SG00-PT. These results serve as primary reference points to ensure fair, environment-matched comparisons. [^1] We report reproduced results rather than official numbers, as the authors of LMDrive acknowledge that performance varies significantly depending on GPU model as reported by other authors as well (see issues 37, 52, and 56 in LMDrive repository; links omitted for review according to author guidelines). Same behaviour was observed for BEVDriver. Additionally, to isolate the effect of scene graph supervision itself, we fine-tune both models on our dataset without scene graph input at training or test-time (SG00-FT). Tab. 1 shows that SG00 fine-tuning shifts performance in a model-specific way: LMDrive gains in safety (higher IS) but sees a drop in RC, while BEVDriver substantially improves RC and resulting in a higher DS. These outcomes provide a neutral reference for assessing the added value of scene graph conditioning.
@@
Specifically, we evaluate all combinations of abstraction level, serialization format, and injection template. Tab. 2 reports DS, RC, and IS for LMDrive averaged over the LangAuto Tiny, Short, and Long tracks. We give mean rows and columns to support comparisons across serialization formats and injection templates, respectively. Notably, multiple test-time scene graph configurations already surpass the reproduced baseline without scene graphs (Tab. 1), underscoring the potential of this approach even without additional training. Two consistent trends emerge: First, injection template V3 outperforms both V2 and V1, with V1 trailing significantly. Second, Actor-Only abstractions yield the best results across all formats, indicating that minimal yet structured relational information is most beneficial when applied to pretrained models.
These results highlight the strength of lean representations and well-designed prompts, even without additional training. Actor-Only graphs in particular strike a favorable balance between semantic relevance and token budget, introducing the fewest additional tokens across all formats, as shown in Tab. 3. Based on these insights, we select the Actor-Only abstraction and injection templates V2 and V3 for all subsequent fine-tuning experiments, focusing on configurations that maximize prompt efficiency and clarity.
@@
Serialization Configuration LMDrive BEVDriver
@@
Scene Graph-Conditioned Fine-Tuning. Tab. 4 reports results for SG11 and SG10 configurations, using the best-performing abstraction (Actor-Only) and injection templates (V2 and V3). For LMDrive, SG10 consistently outperforms SG11 in DS, and V3 surpasses V2, suggesting that exposure to structured context during training enhances internal planning behavior, even when that structure is absent at test-time. The best configuration (Text, Actor-Only, V3, SG10) achieves a DS of 51.8, marking a substantial improvement over the reproduced baseline (42.95) and SG00 fine-tuning (41.55). Moreover, SG10 and SG11 significantly outperform SG01 (44.7), demonstrating that scene graph-conditioned training is more effective than test-time injection only. BEVDriver shows complementary preferences: YAML slightly outperforms Text, and V2 is favored over V3. Nevertheless, SG10 again matches or exceeds SG11, with the top configuration (YAML, Actor-Only, V2, SG10) reaching a DS of 56.1, significantly outperforming both the reproduced baseline (44.70) and SG00 (48.94). We attribute differences in optimal serialization and templates in part to the underlying language backbone (LLaVA-v1.5 in LMDrive vs. LLaMA-7B in BEVDriver), but the broader pattern holds across both: SG10 provides consistent benefits without requiring graph input at test-time.
@@
Extended Infraction Analysis. To better understand how scene graph conditioning influences safety and behavior, we analyze detailed sub-metrics in Tab. 5. Rather than focusing on best-case results, we report averages over all V3-based LMDrive configurations (Text, JSON, YAML) to assess the generality of the observed effects.
@@ 5 Conclusion
[…]
Supplementary Material
This supplementary document provides additional information to support the results presented in the main paper. We include detailed implementation settings, full training-time measurements, extended quantitative results, and additional qualitative analyses. All supplementary content is intended to increase reproducibility and offer full transparency of our experimental pipeline.
6 Implementation Details
This section provides additional details on model configuration, dataset preprocessing, and training hyperparameters that apply to both LMDrive and BEVDriver. These settings remain fixed across all experiments unless explicitly mentioned. Tab. 6 summarizes the shared configuration.
For training and testing, we use eight GPU nodes of a high-performance computing cluster, where each node is equipped with an AMD EPYC 9254 CPU, 384 GB RAM, and an NVIDIA L40s GPU with 48 GB VRAM.
Table 6: Shared Implementation Details for LMDrive and BEVDriver. Both methods use their best-performing language backbones with official pretrained checkpoints (LMDrive: LLaVA-v1.5, BEVDriver: LLaMA-7B). All model, dataset, and training specific hyperparameters are shared across experiments.
Category Hyperparameter
Model freeze_vit: True
use_notice_prompt: False
LoRA rank: 16
LoRA alpha: 32
LoRA dropout: 0.05
Dataset enable_start_frame_augment: True
token_max_length: 40
enable_notice: False
Training Optimizer: AdamW
lr_sched: linear_warmup_cosine_lr
init_lr: 1e-4
min_lr: 1e-5
warmup_lr: 1e-6
warmup_steps: 2000
weight_decay: 0.06
max_epoch: 5
batch_size_train: 1
world_size: 8
7 Training Time
[…]
Serialization Template Full Road-Level Actor-Only
[…]
Serialization Template Full Road-Level Actor-Only
[…]
Table 9: Computational analysis. Measured latency (in milliseconds) and GPU memory usage (in GB) for each stage of the scene graph pipeline (generation, abstraction, serialization) and model inference across abstraction levels (Full, Road-Level, Actor-Only) and serialization formats (Text, JSON, YAML) under the SG11 system configuration using template V3 and LMDrive as base model. For each measurement, we report the 1st percentile (P01), median (Med), and 99th percentile (P99). The final row (SG10) serves as a baseline where no scene graphs are used at test-time, and thus no scene graph pipeline or additional token overhead is involved.
Serialization Abstraction Generation [ms] Abstraction [ms] Serialization [ms] Inference Time [ms] GPU Usage [GB]
P01 Med P99 P01 Med P99 P01 Med P99 P01 Med P99 P01 Med P99
JSON Full 25.0 40.7 53.6 0.0 0.0 0.0 0.3 1.2 2.3 158.9 1259.2 2057.6 25.5 39.6 41.7
Road-Level 25.8 38.6 52.8 0.1 0.2 1.1 0.4 0.9 1.9 203.9 1065.8 1967.4 23.7 36.0 41.9
Actor-Only 25.2 40.1 54.6 0.1 0.1 1.0 0.2 0.5 0.8 185.1 253.9 601.3 25.0 25.1 25.2
YAML Full 25.2 40.9 54.3 0.0 0.0 0.0 1.0 5.7 7.7 225.1 1056.6 1566.7 26.5 32.6 40.1
Road-Level 25.7 41.4 54.7 0.1 0.2 0.9 1.1 4.8 7.5 287.4 924.5 1598.3 27.8 29.5 38.1
Actor-Only 25.6 39.2 54.6 0.1 0.2 1.0 0.8 1.3 2.6 288.3 351.1 516.4 23.1 26.2 29.3
Text Full 26.0 40.6 53.4 0.0 0.0 0.0 0.4 0.9 2.0 79.5 107.4 136.9 20.8 20.9 21.0
Road-Level 25.3 41.4 54.0 0.1 0.2 0.4 0.2 0.8 1.6 58.1 93.3 132.3 19.6 19.6 19.8
Actor-Only 25.4 38.7 53.6 0.1 0.1 0.3 0.2 0.5 0.7 46.4 85.4 118.9 18.2 18.3 18.4
SG10 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 51.6 56.9 65.5 18.2 18.3 18.4
Table 10: Extended baseline performance. Performance of LMDrive and BEVDriver without scene graph input during training and test-time (SG00) on different LangAuto benchmarks Tiny, Short, and Long. SG00-PT refers to the official pretrained checkpoints, while SG00-FT applies further fine-tuning on our collected dataset.
Method Config Tiny Short Long
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
LMDrive SG00-PT 60.72 70.98 0.82 41.34 56.98 0.79 26.80 36.47 0.77
SG00-FT 57.98 65.83 0.88 41.69 50.25 0.86 24.98 28.58 0.90
BEVDriver SG00-PT 63.15 68.31 0.92 42.22 46.00 0.92 28.74 34.79 0.86
SG00-FT 55.80 62.22 0.88 62.27 69.35 0.89 28.75 37.53 0.82
[…]
Across both models, the graph abstraction level is the dominant factor affecting training time, primarily due to differences in sequence length (see Tab. 3 for mean and maximum token counts). The choice of template version has only a negligible impact, and differences between the Text and JSON/YAML serializations follow a consistent pattern: Text yields the shortest training times, while JSON and YAML are comparable but slower due to their higher structural verbosity. The same trends hold for BEVDriver, although its overall training times are substantially shorter than those of LMDrive, which can be attributed to the usage of a different language backbone.
8 Computational Analysis
All experiments in this section were executed on a single NVIDIA L40S GPU node (see Sec. 6). To quantify the computational overhead introduced by scene graph processing, we measure the runtime and resource usage of each component: generation, abstraction, and serialization. These steps are evaluated across all graph abstraction levels (Full, Road-Level, Actor-Only) and serialization formats (Text, JSON, YAML) using template version V3. In addition, since test-time scene graph injection (SG01/SG11) introduces further computational costs, we also report inference latency and GPU memory usage for each configuration with LMDrive as base model.
Tab. 9 presents detailed latency and memory statistics for various SG11 system configurations. Scene graph generation time remains consistent, while abstraction is only applied in Road-Level and Actor-Only configurations, contributing minimally to overall latency. Serialization time varies more significantly: Full and Road-Level abstractions require more tokens and structural elements, especially in JSON and YAML, while Actor-Only remains lightweight.
Inference time is strongly influenced by both abstraction level and serialization format: Text is the most efficient, while JSON and YAML incur higher latency, particularly for Full and Road-Level abstractions due to their verbosity. This trend is also reflected in GPU memory usage. For comparison, we include SG10 system configuration, which excludes scene graph input at test-time, resulting in the lowest overall latency and GPU usage.
9 Detailed Quantitative Results
This section provides full numerical results for all configurations evaluated in the paper. The main paper reports averaged or representative values due to space constraints; here, we include all combinations for completeness.
Extended Baseline Performance. Tab. 10 reports baseline results for LMDrive and BEVDriver without scene graph input (SG00), comparing pretrained (SG00-PT) and fine-tuned (SG00-FT) models across LangAuto Tiny, Short, and Long tracks. This table complements Tab. 1 and provides full benchmark breakdowns for completeness.
Table 11: Extended test-time-only scene graph injection (SG01). Performance across serialization formats, graph abstraction levels, and template versions on different LangAuto benchmarks Tiny, Short, and Long for LMDrive using SG01 system configuration.
Serialization Abstraction Template Tiny Short Long
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
Text Full V1 49.6 56.0 0.87 44.7 54.1 0.85 22.5 28.5 0.81
V2 59.3 64.4 0.89 42.1 52.8 0.83 23.9 31.0 0.77
V3 58.5 66.5 0.86 40.0 49.5 0.84 25.3 33.1 0.79
Road-Level V1 54.2 60.8 0.86 43.8 52.0 0.86 23.8 28.3 0.84
V2 58.2 64.1 0.89 43.5 54.1 0.82 22.9 30.0 0.77
V3 58.3 64.1 0.87 41.1 56.8 0.80 22.1 29.3 0.75
Actor-Only V1 50.9 60.1 0.85 42.9 52.4 0.86 26.3 33.7 0.83
V2 64.7 71.3 0.89 44.8 53.1 0.87 24.5 34.5 0.77
V3 59.8 65.7 0.89 45.3 56.4 0.83 24.2 31.4 0.80
JSON Full V1 45.8 51.4 0.87 34.6 46.3 0.81 23.4 28.8 0.83
V2 61.5 66.9 0.90 39.2 49.7 0.84 23.7 29.4 0.79
V3 62.2 69.3 0.88 42.0 55.0 0.80 29.7 39.6 0.80
Road-Level V1 44.2 49.3 0.88 38.1 54.0 0.78 21.0 26.8 0.82
V2 59.4 66.6 0.86 41.7 56.8 0.77 24.3 30.1 0.78
V3 62.7 71.3 0.86 49.5 59.8 0.85 27.7 37.9 0.75
Actor-Only V1 55.6 60.8 0.89 41.5 51.0 0.86 23.9 30.1 0.81
V2 60.2 64.7 0.91 42.7 58.1 0.78 25.0 34.1 0.76
V3 58.7 71.0 0.82 48.0 58.1 0.86 27.5 37.5 0.77
YAML Full V1 49.4 54.2 0.87 47.0 54.1 0.88 21.7 28.1 0.81
V2 60.5 66.0 0.90 41.0 54.7 0.80 23.9 30.7 0.76
V3 56.7 62.7 0.87 43.5 58.7 0.79 25.2 34.2 0.74
Road-Level V1 52.3 58.0 0.87 43.1 53.8 0.86 22.3 28.0 0.82
V2 59.3 64.8 0.90 41.4 58.3 0.77 24.8 31.6 0.79
V3 57.9 63.8 0.88 43.5 54.4 0.84 24.9 33.7 0.77
Actor-Only V1 60.9 69.3 0.86 43.2 56.2 0.82 24.3 30.6 0.82
V2 60.6 69.3 0.88 44.8 54.0 0.85 25.2 36.1 0.72
V3 62.1 67.5 0.91 46.7 58.1 0.84 25.3 32.6 0.78
Extended Test-Time-Only Scene Graph Injection (SG01). Tab. 11 lists complete SG01 results for LMDrive across all serialization formats, abstraction levels, and prompt templates, broken down by LangAuto benchmark track (Tiny, Short, Long). The results complement Tab. 2.
Extended Scene Graph-Conditioned Fine-Tuning. Tab. 12 and Tab. 13 complement Tab. 4 from the main paper by reporting complete benchmark-wise results for LMDrive and BEVDriver, respectively. All configurations use Actor-Only graphs and compare SG11 (train and test-time) and SG01 (test-time-only) settings across serialization formats and prompt templates V2 and V3. While the main analysis focuses on overall means, these detailed results are provided for completeness.
10 Qualitative Results
Alongside this supplementary material, we include a video showcasing qualitative results of our approach, as well as failure cases of the LMDrive and BEVDriver baselines in which our method demonstrates superior driving performance. The failure scenarios include route deviations caused by incorrect instruction following, lane departure due to poor lane tracking, collisions resulting from limited spatial awareness of other traffic participants, and red light violations stemming from failure to recognize traffic signals. Additionally, the video features examples illustrating different abstraction levels of the scene graph (Full, Road-Level, and Actor-Only) for the same driving scenario.
Table 12: Extended scene graph-conditioned fine-tuning of LMDrive. Actor-Only scene graphs with SG11 and SG01 system configuration across LangAuto benchmarks (Tiny, Short, and Long) for LMDrive using template versions V2 and V3.
Serialization Template Configuration Tiny Short Long
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
Text V2 SG11 66.6 76.1 0.88 45.5 60.9 0.77 27.2 36.9 0.76
SG01 66.8 75.2 0.88 46.4 57.8 0.82 27.3 35.4 0.81
V3 SG11 52.8 62.0 0.85 56.3 65.8 0.87 32.0 40.5 0.78
SG01 72.6 78.7 0.91 52.9 60.6 0.87 29.9 39.0 0.81
JSON V2 SG11 55.2 67.2 0.82 45.9 57.7 0.81 30.1 40.2 0.77
SG01 57.7 67.6 0.86 53.3 59.7 0.88 26.1 33.2 0.81
V3 SG11 64.9 68.6 0.94 51.1 59.1 0.84 30.5 39.0 0.82
SG01 53.5 60.2 0.89 53.9 62.9 0.85 29.2 38.8 0.81
YAML V2 SG11 61.3 67.1 0.90 47.5 61.2 0.79 27.8 38.8 0.75
SG01 63.8 72.1 0.88 49.1 58.1 0.83 30.9 42.1 0.79
V3 SG11 61.9 71.4 0.86 43.7 53.7 0.79 28.6 38.8 0.77
SG01 63.7 71.0 0.89 50.9 56.8 0.87 25.8 36.8 0.75
Table 13: Extended scene graph-conditioned fine-tuning of BEVDriver. Actor-Only scene graphs with SG11 and SG01 system configuration across LangAuto benchmarks (Tiny, Short, and Long) for BEVDriver using template versions V2 and V3.
Serialization Template Configuration Tiny Short Long
DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$ DS $\uparrow$ RC $\uparrow$ IS $\uparrow$
Text V2 SG11 60.7 69.0 0.88 59.2 71.5 0.83 35.2 43.7 0.81
SG01 66.9 78.5 0.85 50.0 54.5 0.93 38.2 46.5 0.83
V3 SG11 61.5 71.1 0.88 60.8 70.4 0.88 34.2 49.0 0.73
SG01 61.3 67.8 0.90 59.5 67.8 0.89 33.7 44.4 0.78
JSON V2 SG11 60.9 69.4 0.88 63.9 73.3 0.88 34.7 45.7 0.76
SG01 52.0 57.1 0.89 54.9 58.5 0.94 35.1 40.7 0.85
V3 SG11 49.3 56.3 0.89 47.2 55.0 0.90 34.9 43.0 0.83
SG01 49.2 56.9 0.90 48.4 55.9 0.89 29.1 39.3 0.77
YAML V2 SG11 58.0 64.0 0.91 64.1 72.8 0.87 36.1 48.1 0.78
SG01 58.8 69.2 0.86 71.7 76.2 0.94 37.7 46.8 0.81
V3 SG11 58.4 67.8 0.86 58.5 63.6 0.90 41.7 49.7 0.84
SG01 58.8 66.4 0.86 55.3 62.9 0.87 30.5 37.7 0.81